Derivatives:Problem 1: On January 2, 2017, Jones Company purchases a call option for $300 on Merchant common stock. The call option gives
Answer #1

Ans: Journal entries

S.no Account title and explanation Debit($) Credit($)
1 Call option   $300
Cash $300
(being call option purchased)
2 Unrealized gain or loss -income $100
Call option(300-200) $100
Call option 3000
Unrealized gain or loss- income($53-50)=3*1000 3000
( to record change in fair value)
3 effect of net income of entering into the derivative transaction
Unrealized holding gain(3000-100) =$2900